Solve log2(x+2) = log2^3x.
Answer by stanbon(48535)

Solve log2(x+2) = log2^3x.
It's hard to tell from your posting whether "2"
is the base or not. I'm going to assume "2" is
not the base: that 10 is the base.
If that is the case, taking the anti-log you get:
2(x+2)=2^(3x)
Graphing the two sides and finding the intersection
I get x=1.2306008...
Cheers,
Stan H.